# A Watermark for Low-entropy and Unbiased Generation in Large Language Models

### v2

@@ 4.1.1 Statistical Test Guarantees of STA-1
Detecting the STA-1 generated text. The detection of STA-1 compares the empirical proportion of green list tokens in the given text against the green list proportion $\gamma$ (Kirchenbauer et al., 2023a). We employ the $z$-test where the null hypothesis ($H_{0}$) is that the text is generated without knowing the green-red list partition. Denote $|S|_{G}$ as the number of green list tokens in this text. Under $H_{0}$, $|S|_{G}$ follows a Bernoulli distribution $B(T,\gamma)$ with a mean of $\gamma T$ and a variance of $\gamma(1-\gamma)T$. The $z$-score is calculated with the empirical $|S|_{G}$ as
$z=\frac{|S|_{G}-\gamma T}{\sqrt{\gamma(1-\gamma)T}}.$ (2)
The alternative hypothesis ($H_{a}$) is that the text is generated with STA-1. Under $H_{a}$, $|S|_{G}$ is expected to be larger than $\gamma T$. We can detect watermarked texts with a certain confidence level if the $z$-score exceeds a $z$ threshold. For example, if $z>2$, we are more than 97.7% confident that the text is watermarked under the one-tail test.
@@
To ensure the effectiveness of the $z$-test, under $H_{a}$, a lower bound on the expectation of $|S|_{G}$ and an upper bound on the variance of $|S|_{G}$ are required. We establish the necessary lower and upper bounds in the following theorem. Because both bounds are related to the Gini index of the LLM output distribution, we define the Gini index first.
Definition 2 (Gini index). Given a discrete probability distribution $p=(p_{1},p_{2},\cdots,p_{N})$, the Gini index of $p$ is defined as
$Gini(p)=\sum_{i=1}^{N}p_{i}(1-p_{i}).$ (3)
A low Gini index implies less uncertainty in the probability distribution, resulting in a low-entropy scenario. Next, we propose the mean and variance bounds of $|S|_{G}$.
@@
Theorem 2. For STA-1 generated text sequences with $T$ tokens, let the random green list have a fixed size of $\gamma|\mathcal{V}|$, and $p^{t}_{i}$ denote the LLM’s raw output probability of the $i$-th token in $\mathcal{V}$ at step $t$, $i=1,2,\cdots,|\mathcal{V}|$, $p^{t}=(p^{t}_{1},p^{t}_{2},\cdots,p^{t}_{|\mathcal{V}|})$. If an STA-1 generated sequence $S$ has an average Gini index larger than or equal to $Gini^{*}$, that is,
$\frac{1}{T}\sum_{t=1}^{T}Gini(p^{t})=\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{|\mathcal{V}|}p^{t}_{i}(1-p^{t}_{i})\geq Gini^{*}.$
Then the expectation of $|S|_{G}$ is at least
@@
$\mathbb{V}(|S|_{G})\leq T[\gamma+(1-\gamma)\gamma Gini^{*}][1-\gamma-(1-\gamma)\gamma Gini^{*}].$ (5)
@@
Example 1. We show an example for a typical $\gamma$. Let $\gamma=0.5$, this bound becomes
$\mathbb{E}(|S|_{G})\geq\frac{1}{2}T+\frac{1}{4}TGini^{*},$ (6)
$\mathbb{V}(|S|_{G})\leq T[\frac{1}{2}+\frac{1}{4}Gini^{*}][\frac{1}{2}-\frac{1}{4}Gini^{*}]=T[\frac{1}{4}-\frac{1}{16}Gini^{*2}].$ (7)
Note that $Gini^{*}$ is the average Gini index. When the generation becomes more uncertain, $Gini^{*}$ increases and we can expect a higher number of green list tokens with a lower variance. Practically in low-entropy scenarios, with probability masses concentrated on one or a few tokens, those tokens are likely to be generated frequently regardless of the green and red list partition in STA-1. Thus, fewer tokens in the green list are expected. This weakens the strength of watermarking methods and makes watermark detection challenging, which is consistent with the theorem.
@@
Corollary 1. Given that Theorem 2 holds, if $Gini^{*}>\tilde{z}/\sqrt{\gamma(1-\gamma)T}$, we have the type II error
$P\left(\frac{|S|_{G}-\gamma T}{\sqrt{\gamma(1-\gamma)T}}\leq\tilde{z}\bigg{|}H_{a}\right)\leq\frac{\overline{\mathbb{V}}}{\overline{\mathbb{V}}+(\underline{\mathbb{E}}-\gamma T-\tilde{z}\sqrt{\gamma(1-\gamma)T})^{2}},$ (8)
@@
The additional condition requires that $Gini^{*}$ must not be excessively low given the threshold value $\tilde{z}$. A higher $Gini^{*}$ increases $\underline{\mathbb{E}}$ and decreases $\overline{\mathbb{V}}$, resulting in a reduced upper bound on the type II error. Therefore, the test has higher statistical power in high-entropy scenarios.
4.1.2 Discussion on the Low-entropy Protocol
Previous work claims that unbiased watermarks can avoid the tradeoff between watermark strength and text quality (Hu et al., 2024). We challenge this claim by first considering the following example.
Example 2. Assuming that the token set only includes two tokens $\mathcal{V}=\{A,B\}$, at a typical step, an LLM outputs the probability of generating $A$ ($p_{A}$) and $B$ ($p_{B}$) as $(p_{A},p_{B})=(0.8,0.2)$. Consider the following two unbiased watermarks. $W_{1}$: with a probability of 0.2 always generating $B$ and with a probability of 0.8 always generating $A$; $W_{2}$: with a probability of 0.5, the probability distribution becomes $(p_{A},p_{B})=(0.9,0.1)$ and with the other probability of 0.5, becomes $(p_{A},p_{B})=(0.7,0.3)$.
In Example 2, one can view the prompt as ‘The quick brown fox jumps over the lazy’, $A$ as the token ‘dog’, and $B$ as all other tokens. It is easy to show that watermarks $W_{1}$ and $W_{2}$ are both unbiased. However, risk-averse people (Pratt, 1978) will prefer watermark $W_{2}$ because $W_{2}$ does not have a possibility that only $B$ will be sampled. $B$ represents unsatisfactory outputs in low-entropy scenarios which could significantly harm text quality, and we want the risk of sampling $B$ to be as low as possible. We refer readers to Appendix D for a conventional example in finance and a better understanding of the analysis via utility theory. At any generation step, let $x_{max}$ denote the token with the maximum probability $p_{max}$. We measure the risk by the variance (Sharpe, 1998) of $p_{max}^{w,k}$ among watermark keys, where $p_{max}^{w,k}$ denotes the altered value of $p_{max}$ with a watermarking method and a key $k$. We show that STA-1 has a lower risk compared to previous unbiased watermarks in the following theorem.
Theorem 3. Assume $11-\alpha\leq p_{max}<$ 1, where $\alpha$ represents the partition hyperparameter used in Dipmark. For the low-entropy protocol in Section 3, the STA-1 method has a lower variance in the probability of generating $x_{max}$ compared to other unbiased methods (including Dipmark, $\gamma$-reweight, and RDW) (Hu et al., 2024; Wu et al., 2024; Kuditipudi et al., 2023). Formally,
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]<\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]=\mathbb{V}_{k\sim P_{K}(k)}^{\gamma\text{-reweight}}\left[p_{max}^{w,k}\right]<\mathbb{V}_{k\sim P_{K}(k)}^{\text{RDW}}\left[p_{max}^{w,k}\right],$ (9)
for any $\alpha\in[0,0.5]$ used in Dipmark, where $p_{max}^{w,k}$ denotes the adjusted probability of the token $x_{max}$ under the respective watermarking method with a key $k\in K$.
Proof. See Appendix C.4.
Example 3. We show a numerical example with $p_{max}=0.8$. For STA-1, based on the proof of the theorem, if the proportion of the green list is 0.5, $\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]=0.0064$. For Dipmark ($\alpha\in[0.2,0.5]$) and $\gamma$-reweight, the variance is $\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]=\mathbb{V}_{k\sim P_{K}(k)}^{\gamma\text{-reweight}}\left[p_{max}^{w,k}\right]=\frac{1}{75}\approx 0.013$. For RDW, the variance is $\mathbb{V}_{k\sim P_{K}(k)}^{\text{RDW}}\left[p_{max}^{w,k}\right]=0.16$.
4.2 Sampling M Then Accepting
@@ 5 Experiments
In this section, we conducted computational experiments to evaluate the performance of STA-1 and STA-M using two public datasets. We benchmarked our methods against various watermarking baselines on text quality and watermark strength. Moreover, we discussed the variance of generation in the low-entropy dataset. Finally, we conducted a robustness analysis of STA against different watermarking attacks.
5.1 Experimental Setup
Datasets and metrics. We employed two public datasets: C4 subset (Raffel et al., 2020; Kirchenbauer et al., 2023a) for news-like text generation and HumanEval (Chen et al., 2021) for code generation. We evaluated the performance of different watermarking methods on text quality and watermark strength. For watermark strength, we set the $z$ threshold as 2 and 2.5 and report the F1-score and AUC of watermark detection. For text quality, we measured perplexity (PPL) and coherence (Gao et al., 2021) for generations on C4; We computed PPL and pass@$k$ scores of code generations (Chen et al., 2021) for HumanEval. We refer readers to Appendix F.1 for more dataset details and the prompt used in each dataset.
Baselines. We chose KGW as the biased watermark baseline (Kirchenbauer et al., 2023a), RDW (Kuditipudi et al., 2023), $\gamma$-reweight (Hu et al., 2024), and Dipmark (Wu et al., 2024) as the unbiased watermark baselines. Specifically, we set KGW with a fixed green list proportion $\gamma=0.5$ and diverse logit increments $\delta\in\{1,1.5,2\}$. We set the watermark key length as 256 in RDW. The partition parameter of Dipmark was set as $\alpha\in\{0.3,0.4,0.5\}$. When $\alpha=0.5$, we report this result as $\gamma$-reweight. Note that $\gamma$-reweight (Hu et al., 2024) does not include a $z$-test. Therefore, we implemented the $z$-score in Dipmark (Wu et al., 2024) for $\gamma$-reweight by counting the number of tokens in the latter portion of the token set. We also show results without watermarking techniques. Also, RDW only contains a permutation test that reports p-values. We set p-value thresholds at 0.05 and 0.01 to approximate two $z$-tests.
Implementation details. We utilized different variants of LLaMA-2-7B (Touvron et al., 2023) as our generative models, and LLaMA-2-13B to compute perplexity. For hyperparameters in STA-M, we set $M\in\{4,8,16\}$ and two entropy thresholds $\tau$ for different datasets. We conducted a robustness check on $\tau$ in Appendix F.2 and selected different $\tau$s for different datasets in the final experiment. For each method, we run 10 times to conduct all pair-wise Tukey tests. Results in the following tables show only average values. We refer readers to Appendix F.1 for more details on implementation.
Table 1: Result Comparison between Our Methods and Baselines on Text Quality and Watermark Strength for the C4 Dataset. The best results without statistical differences are shown in bold. The second best results without statistical differences are shown in underline.
@@ 5.1 Experimental Setup
$z=2.0$ $z=2.5$ Time
Method PPL Coherence F1 AUC F1 AUC
@@
Table 2: Result Comparison between Our Methods and Baselines on Text Quality and Watermark Strength for the HumanEval Dataset. The best results without statistical differences are shown in bold. The second best results without statistical differences are shown in underline.
@@
Method PPL Pass@$1$ Pass@$5$ Pass@$10$ F1 AUC F1 AUC
@@
5.2 Results on C4
For the C4 dataset, each method generates at least 500 text sequences with at least $200\pm 5$ tokens (Kirchenbauer et al., 2023a). Table 1 demonstrates each method’s text quality, watermark strength, and detection time for 500 generations, and we present generated text examples in Appendix F.3. As depicted in Table 1, the proposed STA-1 method is efficient in detection and achieves comparable perplexity and coherence compared to no watermark generation. The text quality results are consistent with other unbiased watermarks including RDW, Dipmark, and $\gamma$-reweight, showing STA-1 is also unbiased empirically. In terms of watermark strength, STA-M ($M\in\{4,8,16\}$) outperforms all unbiased watermarks and has comparable watermark strength as biased watermarks KGW ($\delta\in\{1.5,2\}$). Overall in the high-entropy generation task, the unbiased STA-1 method is comparable to other unbiased watermarks; The STA-M method can improve the watermark strength by sacrificing minor text quality.
@@ 5.3 Results on HumanEval
In this section, we compare our methods against baselines on the HumanEval dataset. Table 2 presents the perplexity, pass@$k$ scores, and watermark strength for all methods. Since it is better not to control the length of a code during generation, we remove detection time results. First, we focus on the result analysis for all unbiased watermarks. As reported, our STA-1 method achieves similar perplexity and pass@$k$ scores compared to no watermarking and other unbiased watermarking methods.
Table 3: Comparison on the Risk of Unsatisfactory Outputs for Unbiased Watermarks. For space concern, we denote the number of passed problems as PP, the number of passed codes as PC, and the average number of passed codes per passed problem as PC per PP (PC/PP).
Method PPL Variance PP PC PC per PP
RDW 2.202 77 219 2.844
Dipmark($\alpha$=0.3) 1.535 84 233 2.675
Dipmark($\alpha$=0.4) 1.853 84 221 2.631
$\gamma$-reweight 1.722 80 214 2.774
STA-1 1.461 81 254 3.136
@@
Moreover, we examine the risk of unsatisfactory outputs produced by unbiased watermarks for low-entropy generations. Specifically, we compare different unbiased watermarks in terms of four more metrics. We ran 10 times of code generation for each problem using different unbiased watermarking methods with 10 different keys. Table 3 reports the average variance of perplexity among each problem, the number of passed problems (if the problem is solved by any one generation out of 10 runs, it is considered passed), the number of passed codes, and the average number of passed codes among all passed problems. In particular, the STA-1 method demonstrates the lowest variance of perplexity compared to RDW, Dipmark($\alpha$=0.3), Dipmark($\alpha$=0.4), and $\gamma$-reweight with a variance of 1.461 compared to 2.202, 1.535, 1.853, and 1.722, respectively. A lower variance indicates a lower risk among different text generations under different keys. Additionally, we show the average number of passed codes among all passed problems. For example, 3.136 in Table 3 means among all solved problems, an average of 3.136 generated codes are accurate w.r.t. 10 generations by STA-1. We conclude from Table 3 that although Dipmark solves more problems, it fails to provide consistent accurate codes among different generations. Instead, our method outperforms other unbiased watermarks (RDW, Dipmark($\alpha$=0.3), Dipmark($\alpha$=0.4), $\gamma$-reweight) in providing consistency, with an average number of passed codes of 3.136 compared to 2.844, 2.675, 2.631, and 2.774, respectively. In summary, the STA-1 method has a lower risk when generating low-entropy texts, as discussed in Theorem 3.
@@
In terms of watermark strength in Table 2, STA-M ($M\in\{4,8,16\}$) yields higher watermark strength in comparison to all unbiased watermarks while maintaining similar pass scores. The STA-16 method achieves comparable watermark strength against biased watermark KGW($\delta=2$) with an AUC of 0.764 ($z=2$) against 0.795. The text quality is maintained with a pass@$10$ of 0.476, highlighting the efficacy of the heuristics to enhance watermark strength at high-entropy generation steps.
Table 4: Attacking Watermarks for the C4 Dataset.
        Attack Setting No Attack Copy-Paste GPT-3.5 DIPPER-1 DIPPER-2
Method F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC
KGW($\delta=1$) 0.96 0.96 0.68 0.75 0.27 0.57 0.13 0.53 0.15 0.54
KGW($\delta=1.5$) 0.99 0 98 0.90 0.90 0.41 0.62 0.22 0.56 0.27 0.57
KGW($\delta=2$) 0.99 0.99 0.95 0.95 0.54 0.68 0.30 0.58 0.40 0.62
RDW 0.98 0.98 0.83 0.79 0.73 0.78 0.64 0.73 0.65 0.73
Dipmark($\alpha=0.3$) 0.93 0.94 0.61 0.70 0.29 0.57 0.24 0.55 0.26 0.55
Dipmark($\alpha=0.4$) 0.96 0.96 0.75 0.79 0.38 0.61 0.31 0.58 0.34 0.59
$\gamma$-reweight 0.96 0.96 0.74 0.78 0.41 0.61 0.32 0.57 0.36 0.60
STA-1 0.96 0.96 0.78 0.81 0.47 0.63 0.39 0.60 0.46 0.63
STA-4($\tau$=1.35) 0.97 0.97 0.95 0.95 0.72 0.78 0.65 0.73 0.69 0.75
STA-8($\tau$=1.35) 0.98 0.98 0.95 0.95 0.78 0.81 0.71 0.77 0.76 0.79
STA-16($\tau$=1.35) 0.97 0.97 0.95 0.95 0.76 0.80 0.68 0.74 0.78 0.81

### v3

@@ 3.1.2 Statistical Test Guarantees
The proposed STA-1 method also has a statistical test guarantee of type II error for detection. Specifically, the detection of STA-1 compares the empirical proportion of green list tokens in the given text against the green list proportion $\gamma$ (Kirchenbauer et al., 2023a). We employ the $z$-test where the null hypothesis ($H_{0}$) is that the text is generated without knowing the green-red list partition. Denote $|S|_{G}$ as the number of green list tokens in this text. Under $H_{0}$, $|S|_{G}$ follows a Binomial distribution $B(T,\gamma)$ with a mean of $\gamma T$ and a variance of $\gamma(1-\gamma)T$. The $z$-score is calculated with the empirical $|S|_{G}$ as $z=\frac{|S|_{G}-\gamma T}{\sqrt{\gamma(1-\gamma)T}}.$ The alternative hypothesis ($H_{a}$) is that the text is generated with STA-1. Under $H_{a}$, $|S|_{G}$ is expected to be larger than $\gamma T$. We can detect watermarked texts with a certain confidence level if the $z$-score exceeds a $z$ threshold.
@@
To ensure the effectiveness of the $z$-test, under $H_{a}$, a lower bound on the expectation of $|S|_{G}$ and an upper bound on the variance of $|S|_{G}$ are required. We establish the necessary lower and upper bounds in the following theorem. We first define the Gini index as $Gini(p)=\sum_{i=1}^{N}p_{i}(1-p_{i})$, where $p$ is a discrete probability distribution given by $(p_{1},p_{2},\cdots,p_{N})$. A low Gini index implies less uncertainty in the probability distribution, resulting in a low-entropy scenario. Next, we propose the mean and variance bounds of $|S|_{G}$.
@@
Theorem 3. For STA-1 generated text sequences with $T$ tokens, let the random green list have a fixed size of $\gamma|\mathcal{V}|$, and $p^{t}_{i}$ denote the LLM’s raw output probability of the $i$-th token in $\mathcal{V}$ at step $t$, $i=1,2,\cdots,|\mathcal{V}|$, $p^{t}=(p^{t}_{1},p^{t}_{2},\cdots,p^{t}_{|\mathcal{V}|})$. If an STA-1 generated sequence $S$ has an average Gini index larger than or equal to $Gini^{*}$, that is, $\frac{1}{T}\sum_{t=1}^{T}Gini(p^{t})\geq Gini^{*}$, then the expectation of $|S|_{G}$ is at least
@@
$\mathbb{V}(|S|_{G})\leq T$ $[\gamma+(1-\gamma)\gamma Gini^{*}]$ (4)
$[1-\gamma-(1-\gamma)\gamma Gini^{*}].$
@@
Corollary 1. Given that Theorem 3 holds, if $Gini^{*}>\tilde{z}/\sqrt{\gamma(1-\gamma)T}$, we have the type II error $\beta=P(\frac{|S|_{G}-\gamma T}{\sqrt{\gamma(1-\gamma)T}}\leq\tilde{z}\big{|}H_{a})$ satisfy
$\beta$ $\leq\frac{\overline{\mathbb{V}}}{\overline{\mathbb{V}}+(\underline{\mathbb{E}}-\gamma T-\tilde{z}\sqrt{\gamma(1-\gamma)T})^{2}},$ (5)
@@
A higher $Gini^{*}$ increases $\underline{\mathbb{E}}$ and decreases $\overline{\mathbb{V}}$, resulting in a reduced upper bound on the type II error. Therefore, the test has higher statistical power in high-entropy scenarios.
3.2 Sampling M Then Accepting (STA-M)
@@ 4 Experiments
In this section, we conducted computational experiments to evaluate the performance of STA-1 and STA-M using two public datasets. We benchmarked our methods against various watermarking baselines on text quality, watermark strength, and detection time. Moreover, we discussed the risk of unsatisfactory outputs in the low-entropy dataset. Finally, we conducted a robustness analysis of STA-1 and STA-M against different watermarking attacks.
Table 2: Result Comparison between Our Methods and Baselines on Text Quality and Watermark Strength for the C4 Dataset. For unbiased watermarks, the best results without statistical differences are underlined. For biased watermarks, the best results without statistical differences are shown in bold.
@@
$z=2.0$ $z=2.5$ Efficiency
Method $\downarrow$ PPL $\uparrow$ Coherence $\uparrow$ F1 $\uparrow$ AUC $\uparrow$ F1 $\uparrow$ AUC Total Time
@@
4.1 Experimental Setup
Datasets and metrics. We employed two public datasets: C4 subset (Raffel et al., 2020; Kirchenbauer et al., 2023a) for news-like (high-entropy) text generation and HumanEval (Chen et al., 2021) for code (low-entropy) generation. We evaluated the performance of different watermarking methods on text quality and watermark strength. For text quality, we measured perplexity (PPL) and coherence (Gao et al., 2021) for generations on C4; We computed PPL and pass@$k$ scores of code generations (Chen et al., 2021) for HumanEval. We refer readers to Appendix F.1 for more dataset details and the prompt used in each dataset. For watermark strength, we set the $z$ threshold as 2 and 2.5 and report the F1-score and AUC of watermark detection. Additionally, for the C4 subset, we employed true positive rate at false positive rate (TPR@FPR) as a metric to evaluate the detection Liu et al. (2023a).
Baselines. We chose KGW as the biased watermark baseline (Kirchenbauer et al., 2023a), RDW (Kuditipudi et al., 2023), Dipmark (Wu et al., 2024), and $\gamma$-reweight (Hu et al., 2024) as the unbiased watermark baselines. Specifically, we set KGW with a fixed green list proportion $\gamma=0.5$ and diverse logit increments $\delta\in\{1,1.5,2\}$. We set the watermark key length as 256 in RDW. The partition parameter of Dipmark was set as $\alpha\in\{0.3,0.4,0.5\}$. When $\alpha=0.5$, we report this result as $\gamma$-reweight. Note that $\gamma$-reweight (Hu et al., 2024) does not include a $z$-test. Therefore, we implemented the $z$-score in Dipmark (Wu et al., 2024) for $\gamma$-reweight by counting the number of tokens in the latter portion of the token set. Also, RDW only contains a permutation test that reports p-values. We set p-value thresholds at 0.05 and 0.01 to approximate two $z$-tests.
Implementation details. We utilized different variants of LLaMA-2-7B (Touvron et al., 2023) as our generative LLMs, and LLaMA-2-13B to compute perplexity. For hyperparameters in STA-M, we set $M\in\{4,8,16\}$ and two entropy thresholds $\tau$ for different datasets. We conducted a robustness check on $\tau$ in Appendix F.2 and selected different $\tau$s for different datasets in the final experiment. For each method, we run 10 times to conduct all pair-wise statistical tests. Results in the following tables show only average values. For detection efficiency, we report the detection time for all generations.
4.2 Results on C4
Figure 1: Result Comparison of Watermark Strength of TPR@0.1%FPR Between Our Method and Baselines for the C4 Dataset.
For the C4 dataset, each method generates at least 500 text sequences with at least $200\pm 5$ tokens (Kirchenbauer et al., 2023a). Table 2 demonstrates each method’s text quality, watermark strength, and detection efficiency for 500 generations, and we present generated text examples in Appendix F.3. We can observe that the proposed STA-1 method achieves comparable perplexity and coherence when compared to no watermark generation and existing unbiased watermarks, including RDW, Dipmark, and $\gamma$-reweight. This result empirically shows that the STA-1 method is unbiased. In terms of watermark strength, the STA-1 method also achieves satisfactory results on F1 and AUC and is not inferior to existing unbiased benchmarks. We also plot the detection performance of TPR@0.1%FPR in Figure 1. We observe that, unlike some unbiased watermarks such as Dipmark, which experience a significant drop in TPR@0.1%FPR, the STA-1 method remains comparable to the best-performing unbiased watermark RDW. Furthermore, based on Table 2, the STA-1 method is highly efficient, taking only around 40 seconds to detect 500 generations.
In Table 2, we also report the performance of STA-M, which samples more times at high-entropy steps for improving watermark strength. In terms of watermark strength, it is evident that STA-M ($M\in\{4,8,16\}$) outperforms all unbiased watermarks and demonstrates results comparable to the biased KGW watermark ($\delta\in\{1.5,2\}$). We also plot the TPR@0.1%FPR scores for different parameter settings of STA-M and KGW in Figure 1. From this, we observe that the watermark strength of KGW ($\delta\in\{1,1.5,2\}$) varies significantly, ranging from 0.7 to 0.99. In contrast, our STA-M ($M\in\{4,8,16\}$) method remains stable across all parameter settings, with a consistent TPR@0.1%FPR score over 0.99. Regarding text quality, STA-M does not experience significant drops compared to the unbiased watermarks and remains comparable to the biased KGW methods.
Table 3: Result Comparison between Our Methods and Baselines on Text Quality and Watermark Strength for the HumanEval Dataset. For unbiased watermarks, the best results without statistical differences are underlined. For biased watermarks, the best results without statistical differences are shown in bold.
@@ 4.2 Results on C4
Method $\downarrow$ PPL $\uparrow$ Pass@$1$uparrow Pass@$5$uparrow Pass@$10$uparrow F1 $\uparrow$ AUC $\uparrow$ F1 $\uparrow$ AUC
@@ 4.3 Results on HumanEval
We then compare our methods against baselines on the HumanEval dataset, a low-entropy code generation benchmark. We report perplexity, pass@k scores, and watermark strength for all methods in Table 3. Since it is preferable not to control the length of code during generation, we remove detection time results. We observe that our STA-1 method achieves similar perplexity, pass@k scores, and watermark strength compared to other unbiased watermarking methods. This further empirically corroborates that the STA-1 method is unbiased.
@@
Moreover, we examine the risk of unsatisfactory outputs produced by unbiased watermarks for low-entropy generations. Specifically, we ran 10 times of code generation for each problem using different unbiased watermarking methods with 10 different keys. We compute the average variance of perplexity for each problem, as well as the average number of passed codes among all passed problems. The results are shown in Figure 2. From the figure, it is evident that the STA-1 method demonstrates the lowest variance of perplexity compared to RDW, Dipmark($\alpha$=0.3), Dipmark($\alpha$=0.4), and $\gamma$-reweight. A lower variance indicates a lower risk among different text generations under different keys. Additionally, we observe that for STA-1, the average number of passed codes among all passed problems is significantly larger than that of other unbiased watermarks, exceeding 3.1, while others remain below 2.9. Therefore, we can conclude that the STA-1 method has a lower risk when generating low-entropy texts compared to existing unbiased watermarks, as discussed in Theorem 2.
Figure 2: Comparison on the Risk of Unsatisfactory Outputs for Unbiased Watermarks. For space concern, we denote the average number of passed codes among all passed problems as PC per PP.
@@
We then report the performance of STA-M in Table 3. As shown, by repeating sampling during high-entropy steps, STA-M ($M\in{4,8,16}$) achieves higher watermark strength compared to all unbiased watermarks, while maintaining similar pass scores. Specifically, the STA-16 method achieves comparable watermark strength against biased watermark KGW($\delta=2$) with an AUC of 0.764 ($z=2$) against 0.775. Meanwhile, the text quality is maintained with a pass@$10$ of 0.476, highlighting the efficacy of the heuristics to enhance watermark strength at high-entropy generation steps.